OpenStudy (anonymous):

SOLVE THE EQUATION -8X^2+8X-2=0

OpenStudy (anonymous):

\[-8x ^{2} +4x+4x-2\]

OpenStudy (anonymous):

\[-4x*(2x-1)+2*(2x-1)\]

OpenStudy (anonymous):

\[(4x-1)*(-4x+2)\]

OpenStudy (anonymous):

so x= 1/2 or1/4

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

can you help me with -24x2+28/3x-5/6=0

OpenStudy (anonymous):

ok i will give you formula for solving any qudratic equation, if \[ax ^{2} +bx+c=0\] \[x=\frac{ -b+\sqrt{b ^{2}-4ac} }{ 2a }\] or\[\frac{ -b-\sqrt{b ^{2}-4ac} }{ 2a }\]

OpenStudy (anonymous):

okay thanks friend (:

OpenStudy (anonymous):

no problem (: